What is (a) the highest, (b) the lowest total resistance that can be secured by combinations of four coils of resistance 4 Ω , 8 Ω , 12 Ω , 24 Ω?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(a) The highest resistance is when the resistances are connected in series:
R1 = 4 ohmR2 = 8 ohm
R3 = 12 ohmR4 = 24 ohm
Total resistance in series = R1 + R2 + R3 + R4
= 4 + 8 + 12 + 24
= 48 ohm
Thus, highest resistance is 48 ohm.
(b)The lowest resistance is when the resistances are connected in parallel
Total resistance in parallel
= 1/R1 + 1/R2 + 1/R3 + 1/R4
1/R = 1/2 + 1/8 + 1/12 + 1/24
= 12/24
1/R = 1/2 ohm, R = 2 ohm
Thus, lowest resistance is 2 ohm.

Thanks